# SMITIN: Self-Monitored Inference-Time INtervention for Generative Music Transformers

### v1

@@ 4.4 Soft-weighting
This ablation study justifies the soft-weighting technique over selecting top-$K$ heads for ITI. The rationale for head-weighting arises from the observation that the optimal set of top-$K$ heads varies depending on the task or instrument. By adopting soft-weighting, we can bypass the need for hyper-parameter tuning specific to each task, simplifying the ITI process.
@@
To validate this approach, we compare the performance of the full SMITIN (with soft-weigthing) against SMITIN where the initial weights $w_{l,h}(t_{0})$ are instead 1 for the top-$K$ most accurate heads according to probing (i.e., $(l,h)\in\mathcal{H}_{K}$), and 0 otherwise. Results from Table 4.1 indicate that while top-$K$ selection can yield good performance, it often requires task-specific tuning of $K$. In contrast, soft-weighting demonstrates a well-balanced performance across various metrics without necessitating such tuning. This balance is crucial because it means soft-weighting can adapt to various musical tasks and preferences, making ITI more flexible and user-friendly. However, it is important to acknowledge that soft-weighting may not always be the optimal choice for every ITI application. Users might prefer the top-$K$ approach depending on their specific needs and the nature of their musical goals.
5 Conclusions and Future Work
We proposed a novel approach for inference-time control of generative music transformers, which self-monitors probe accuracy to impose desired musical traits while maintaining overall music quality. A limitation of our approach is the dependence of performance on probe accuracy, that is if the underlying pre-trained transformer has not accurately learned a concept, or if the probe training set is inadequate, the success rate of our intervention technique may suffer (although it still could be creatively useful). In the future, we plan to explore additional intervention using additional musical traits such as genre, emotion, etc. Furthermore, we hope to investigate building “knobs” for interactively controlling the hyper-parameters necessary for inference-time control, in an effort to enable new technical and creative interactions.
Appendix A Subjective Listening Test Setup and Detailed Results
In selecting files for the listening test discussed in Section 3.5, for each algorithm and instrument addition intervention, we randomly select 20 files above the median objective success rate, and 20 files below the median. This leads to 1280 generated files to be rated (20 files x 2 top/bottom x 4 instruments x 4 algorithms x 2 audio continuation/text-to music). We additionally include real music samples from MoisesDB as references. All audio files included in the listening test were ten seconds in length and normalized to -12 loudness units full scale (LUFS) (Grimm et al., 2010). We obtain at least 3 ratings per file. For the audio continuation experiments, we ask raters to ignore the first 3 seconds of each audio file, as that conditioning signal intentionally does not include the target instrument.
In addition to overall quality, we also ask whether listeners can detect the presence of the instruments targeted by our intervention. This serves to help validate our success rate objective metric. Table 8 compares the Spearman correlation coefficient $\rho$ computed between the success rate objective metric and average human rating computed across all processing algorithms. All correlation coefficients in Table 8 have p-value less than 4e-5. We also compare the average human score with the average success rate objective metric (Alg.) for the top-ranked and bottom-ranked files selected for the listening test. For the top ranked files the scores appear to match quite well, while for the bottom ranked files, it seems human raters tend to overestimate the presence of most instruments compared to the objective metric. We also hypothesize that the lower match for audio continuation in Table 8 compared to text to music may be due in part to the test being more subjectively difficult as raters have to focus on the end of the audio file under test.
Method Audio Continuation Text-to-music
drums bass guitar piano avg drums bass guitar piano avg
MoisesDB 3.98 $\pm$ 0.29
Table 8: Subjective results ranking presence of instruments added by intervention, across multiple algorithms.
Audio Continuation Text-to-music
Top Ranked Bottom Ranked Top Ranked Bottom Ranked
Instrument $\rho$ Human Alg. Human Alg. $\rho$ Human Alg. Human Alg.
drums $0.76$ $0.88$ $0.81$ $0.19$ $0.13$ $0.80$ $0.95$ $0.95$ $0.16$ $0.18$
bass $0.42$ $0.78$ $0.89$ $0.48$ $0.06$ $0.58$ $0.80$ $0.96$ $0.38$ $0.13$
guitar $0.40$ $0.78$ $0.76$ $0.52$ $0.15$ $0.67$ $0.86$ $0.88$ $0.34$ $0.14$
piano $0.32$ $0.60$ $0.72$ $0.33$ $0.32$ $0.55$ $0.70$ $0.79$ $0.31$ $0.23$
avg. $0.48$ $0.76$ $0.80$ $0.39$ $0.17$ $0.67$ $0.83$ $0.90$ $0.30$ $0.17$
Appendix B Objective Evaluation
B.1 Instrument Removal
We explore an additional application: removing the target instrument during Audio Continuation using the same probes employed for instrument addition. To accomplish this, we apply negative intervention strength $\alpha$ to eliminate the target instrument from given input music. The hyperparameter settings for this task are configured as $\alpha=-10.0$ for both original ITI and SMITIN, with $s=1$ and $\tau=0.5$ for SMITIN. The success rate in this context is defined as $\bar{\mathcal{C}}(t)<\tau$, where $\tau=0.5$.
Table B.1 reveals an inverse tendency for each instrument’s removal: drums and bass tend to be continuously generated, whereas guitar and piano are more likely to be omitted during the generation process. Moreover, we find that both text-conditioning prompts: “remove <inst.>”, and “no <inst.>” are ineffective in removing the target instrument. This outcome suggests another aspect of controllability that is challenging to achieve with text prompts alone.
Table 9: Objective evaluation on Instrument Removal.
[…]
B.2 Intervention with Different Directions
Following (Li et al., 2023), we compare intervention results using directions as logistic regression classifiers’ weights $\tilde{\theta}_{l,h}$ and mass mean shift. The intervention direction for mass mean shift is determined during probing as a vector that directs from the centroid of negative activations to the centroid of positive activations. The results are shown in Table 10. Similar to (Li et al., 2023), we observe that using mass mean shift as the intervention direction better achieves the target musical factor (according to success rate). However, we again observe a trade-off between controllability and generation distribution shift. Considering that generation quality is crucial in the music generation task, we adopt $\tilde{\theta}_{l,h}$ as the final intervention direction.
@@ B.2 Intervention with Different Directions
B.3 FAD Compared with Unconditioned Generation
To support the claim that SMITIN does not significantly alter the original distribution of MusicGen, we compute the FAD score using the distribution of unconditioned generation music outputs. For Text-to-Music, the comparison distribution is that of text-conditioned generation using MusicCaps text prompts, as discussed in Table 3.2. According to Table B.3, the FAD score of “add <inst.>” is significantly lower in Text-to-Music than in Audio Continuation, indicating that the output distribution between unconditional and conditional generation experiences a significant shift. We observe that SMITIN with $\alpha=5.0$ exhibits the closest distribution distance to the original MusicGen. This supports the claim that SMITIN achieves the desired controllability while retaining MusicGen’s generation capability.
@@ B.3 FAD Compared with Unconditioned Generation
B.4 Detailed Results of Ablation Studies
We disclose objective results on ablation studies according to each instrument. Table 12 and 13 show the full results of Table 4 (Section 4.1), and Table B.4 shows the full instrument-wise results of Table 4.2 (Section 4.3).
Table 12: Intervention with Multiple Directions - Full Results (1)
Individual Success Rate [%]
start with drums start with bass start with guitar start with piano
Table 13: Intervention with Multiple Directions - Full Results (2)
Table 14: Effects on Number of Probing Data - Full Results (shows full-set success rate)
Number of Paired Data Success Rate* [%] ($\uparrow$) FADL-CLAP ${}_{\text{L-CLAP mus}}$ ($\downarrow$) Similarity ($\uparrow$) Probes Max. Acc [%] ($\uparrow$)
drums bass guitar piano avg. drums bass guitar piano avg. drums bass guitar piano avg. drums bass guitar piano avg.
$n$ = 10 $28.3$ $42.9$ $48.4$ $4.5$ $31.0$ 0.344 $0.267$ $0.445$ $0.373$ $0.357$ $0.855$ $0.928$ $0.920$ $0.935$ $0.909$ $88.6$ $85.3$ $73.3$ $69.0$ $79.1$
$n$ = 100 $25.0$ $48.9$ $45.0$ $8.3$ $31.8$ 0.368 $0.289$ $0.406$ $0.336$ $0.349$ $0.866$ $0.920$ $0.929$ $0.941$ $0.914$ $91.6$ $87.8$ $78.3$ $72.8$ $82.6$
$n$ = 500 $24.5$ $41.0$ $38.2$ $5.8$ $27.3$ 0.360 $0.280$ $0.395$ $0.330$ $0.341$ $0.865$ $0.923$ $0.934$ $0.940$ $0.915$ $93.6$ $88.8$ $81.0$ $74.1$ $84.4$
$n$ = 1000 $25.1$ $37.6$ $34.1$ $5.5$ $25.5$ 0.357 $0.273$ $0.400$ $0.331$ $0.340$ $0.861$ $0.927$ $0.930$ $0.940$ $0.914$ $94.3$ $89.1$ $81.4$ $74.7$ $84.8$
[…]
Appendix C Analytical Insights from Probing
C.1 Exploring a wider range of probing tasks
To expand our understanding of MusicGen’s capabilities beyond recognition on instrument and realism, we conducted probing on a set of more general music understanding tasks. These include music tagging (MTT) (Law et al., 2009), genre classification (GTZAN) (Tzanetakis & Cook, 2002), key detection (GS) (Knees et al., 2015), and emotion recognition (EMO) (Soleymani et al., 2013). We adopt appropriate regression models for probing for these tasks: multinomial logistic regression for multi-class classification tasks, multiple logistic regressions for multi-label classification tasks, and linear regression for regression tasks. For these tasks, we set the input duration as 29.0 seconds, and let the probes take the last time step activation output for classification.
Table 15 presents the probing results of generative music transformers, including various configurations of MusicGen and a comparative model, Jukebox (Dhariwal et al., 2020). We include objective results reported in (Castellon et al., 2021) for Jukebox probing. Jukebox, with its larger number of parameters and activation dimension size, generally shows superior performance. However, our primary interest lies not in the superiority of one model over another but in demonstrating MusicGen’s comprehensive understanding of various musical aspects, indicative of its versatility for ITI.
The insights gained from these probing results reveal MusicGen’s proficiency across a spectrum of musical elements, extending from instrument recognition to more complex attributes like genre and emotion. This opens up possibilities for controlled and targeted music generation, where MusicGen can be guided to produce outputs aligning with specific musical characteristics or themes.
Table 15: Probing generative music transformers.
Dataset MTT GTZAN GS EMO
Task Tagging Genre Key Emotion Num. Param.
[…]
Jukebox $91.5$ $41.4$ $79.7$ $66.7$ $72.1$ $61.7$ $69.9$ 4.8K 5B
[…]
C.2 Full Probing Results
We disclose heatmaps of probing classifier performances on every probing task performed in this paper. Heatmaps from different model configurations are normalized in the same scale for each task.

### v2

@@ 6.5 Soft-weighting
This ablation study justifies the soft-weighting technique over selecting top-$K$ heads for ITI. The rationale for soft-weighting arises from the observation that the optimal set of top $K$ heads varies depending on the task or instrument. By adopting soft-weighting, we can bypass the need for hyper-parameter tuning specific to each task, simplifying the ITI process.
@@
To validate this approach, we compare the performance of the full SMITIN (with soft-weighting) against SMITIN where the initial weights $w_{l,h}(t_{0})$ are instead 1 for the top-$K$ most accurate heads according to probing (i.e., $(l,h)\in\mathcal{H}_{K}$), and 0 otherwise. Results from Table 6.4 indicate that while top-$K$ selection can yield good performance, it often requires task-specific tuning of $K$. In contrast, soft-weighting demonstrates a well-balanced performance across various metrics without necessitating such tuning. This balance is crucial because it means soft-weighting can adapt to various musical tasks and preferences, making ITI more flexible and user-friendly. However, it is important to acknowledge that soft-weighting may not always be the optimal choice for every ITI application. Users might prefer the top-$K$ approach depending on their specific needs and the nature of their musical goals.
@@
6.6 Intervention with Different Directions
Following [27], we compare results using different intervention directions: logistic regression classifiers weights $\tilde{\theta}_{l,h}$ and mass mean shift. The intervention direction for mass mean shift is determined during probing as a vector that directs from the centroid of negative activations to the centroid of positive activations. The results are shown in Table 16. Similar to [27], we observe that using mass mean shift as the intervention direction better achieves the target musical factor (according to success rate). However, we again observe a trade-off between controllability and generation distribution shift. Considering that generation quality is crucial in the music generation task, we adopt $\tilde{\theta}_{l,h}$ as the final intervention direction.
6.7 FAD Compared with Unconditioned Generation
To support the claim that SMITIN does not significantly alter the original distribution of MusicGen, we compute the FAD score using the distribution of music outputs from unconditioned generation. For text-to-music, the comparison distribution is that of text-conditioned generation using MusicCaps text prompts, as discussed in the main paper. According to Table 6.5, the FAD score of “add <inst.>” is significantly lower in text-to-music than in audio continuation, indicating that the output distribution between unconditional and conditional generation experiences a significant shift. We observe that SMITIN with $\alpha=5.0$ exhibits the closest distribution distance to the original MusicGen. This supports the claim that SMITIN achieves the desired controllability while retaining MusicGen’s generation capability.